[Hallc_running] [New Logentry] Hall C RC Update: Monday, July 8

gaskelld at jlab.org gaskelld at jlab.org
Mon Jul 8 09:50:02 EDT 2019

Logentry Text:
Beam setup is progressing after the optics setup restart yesterday.  Beam is not expected in Hall C during DAY shift today or SWING shift - both those shifts are on Standby (target operators should report).  Preliminary estimates suggest beam in Hall C Tuesday on Day shift,  but Tuesday Owl shift should report in case the setup goes faster than expected.

Hall C was taken to Controlled Access and the HMS magnets ramped down for further investigation of the water leak that was found yesterday evening. This morning the source was more precisely identified (appears to be at a juncture between copper pipe and rubber hose near the front of the HMS).   Engineering Installation has been contacted, and will take a look later today.

We will remain in Controlled Access at least until the water leak is resolved.  Those wishing to do work in the hall should contact the RC and then check in with the Counting House before entering the hall.

There will be a brief RC meeting right after the 3 pm Hall C meeting (in A110).  Bluejeans link:  https://bluejeans.com/861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3700182
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190708/58227c5d/attachment.html>

More information about the Hallc_running mailing list